OpenStudy (anonymous):

chain rule!!!! Let f(x) = x^3 + 6 x and g(x) = x^2 + 5. Then g(f(x)) = x^A + B x^C + D x^2 + E they want me to find a,b,c,d,e

OpenStudy (anonymous):

i think i use the chain rule?

OpenStudy (zarkon):

the chain rule is used in calculus...you don't need calculus here

OpenStudy (anonymous):

wait so all i need to do is FOIL?

OpenStudy (zarkon):

do the composition and multiply it out

OpenStudy (anonymous):

what is FOIL? Anw they wnat to find g(f(x)) So which means you will substite x= f(x) into g(x) g(f(x))= (x^3+6x)^2 +5 = x^6 +12 x^4 + 36x^2 +5

OpenStudy (anonymous):

then i just use the chain rule?

OpenStudy (anonymous):

this is applied calculus so it would have to bethe chain rule

OpenStudy (zarkon):

calculus is not needed for this problem.

OpenStudy (anonymous):

I agree, this is just a function problem.

OpenStudy (zarkon):

you might want to look up what the chain rule is